Solve (y ^ 2 * z)/x * p + xzq = y ^ 2 .


1
Expert's answer
2022-02-08T10:59:01-0500

Given "(y^2+z^2)p - xyq + xz = 0"

"\implies (y^2+z^2)p - xyq =- xz"

So, by Lagrange's method, we have

"\frac{dx}{y^2+z^2} = \frac{dy}{-xy} = \frac{dz}{-xz}" ______________(1)

"\implies \frac{x dx + ydy + zdz}{xy^2 + xz^2 - xy^2 - xz^2} = \frac{dz}{-xz}"

"\implies x dx + ydy + zdz=0"

So by integration, we get "x^2 + y^2 + z^2 = c_1" _____________(2)

Also from last two parts of equation (1), we have

"\frac{dy}{-xy} = \frac{dz}{-xz}"

"\implies \frac{dy}{y} = \frac{dz}{z}"

So by integration, we get

"\ln(y) = \ln(z) + \ln(c_2)"

"\implies \frac{y}{z} = c_2" __________(3)

Now, from (2) and (3), solution of given differential equation is

"c_2 = f(c_1) \\
\implies \frac{y}{z} = f(x^2 + y^2 + z^2)" is the solution.